# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ Appendix F Proof of Lemma 4
By the second-order Taylor expansion, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
The gradient of $p_{s,a}(\theta)$ is given by
For the first-order term, we have
$\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})^{\top}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})V(s^{\prime})$
[…]
where in the first inequality, we denote $\mathcal{S}_{h,s,a}^{+}$ as the subset of $\mathcal{S}_{h,s,a}$ such that $\phi_{s,a}^{s^{\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*}-\widetilde{\theta}_{k,h})$ is non-negative, the second inequality holds by the Holder’s inequality, and the third inequality is by the confidence set in Lemma 3.
For the second-order term, we define $u_{s,a}^{s^{\prime}}(\theta)=\phi_{s,a}^{s^{\prime}}\theta$, and $p_{s,a}^{s^{\prime}}(u)=\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}}\exp(u_{s,a}^{s^{\prime\prime}})}$, further define
$F(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})},\widetilde{F}(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})V(s^{\prime})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})}.$
Then, we have
$\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})$
=\ = $\frac{1}{2}\big{(}u(\widehat{\theta}_{k,h})-u(\theta_{h}^{*})\big{)}^{\top}\nabla^{2}\widetilde{F}(u(\bar{\theta}))\big{(}u(\widehat{\theta}_{k,h})-u(\theta_{h}^{*})\big{)}$
=\ = $\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\big{(}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{)}^{\top}\frac{\partial^{2}\widetilde{F}(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{)}$
[…]
where the inequality holds by $V(s)\in[0,H],\forall s$.
$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
$+\frac{3H}{2}\sum_{s^{\prime}}\big{(}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta})).$ (30)
To bound the first term, by applying the AM-GM inequality, we obtain
$H\sum_{s^{\prime}}\sum_{s^{\prime\prime}\neq s^{\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
$\hskip 56.9055pt+\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}$
[…]
Plugging (31) into (30), we have
$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime}}(\theta_{h}^{*})\big{\rvert}\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h})-u_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\big{\rvert}$
[…]
=\ = $\frac{5H}{2}\sum_{s^{\prime}}\big{(}(\phi_{s,a}^{s^{\prime}})^{\top}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta}))$
[…]
where the last inequality holds by Lemma 3. Combining (29) and (32) finishes the proof. ∎
Appendix G Proof of Theorem 2
@@
=\ = $\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(c)}+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(d)}$
$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(e)}.$
@@
=\ = $\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\Big{\lvert}\Big{(}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime}}-p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime\prime}}\Big{)}^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\Big{\rvert}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
@@
$\hskip 28.45274pt+\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\big{\lvert}(\phi_{k,h}^{s^{\prime\prime\prime}})^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\big{\rvert}$
@@
where and the fifth inequality is by Cauchy-Schwarz inequality and the fourth inequality is because by Lemma 17, since ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
$\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}\leq\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\widetilde{\mathcal{H}}_{k,h}}\leq 2\eta\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{\widetilde{\mathcal{H}}_{k,h}^{-1}}\leq\frac{2\eta}{\sqrt{\lambda}}\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2},$
and since $g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
$\lVert g_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2}\leq\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}+\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}\leq 2\max_{s^{\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime}}\big{\rVert}_{2}\leq 2.$
@@
$\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\frac{4\eta}{\sqrt{\lambda}}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}$
@@
where the last inequality holds by Lemma 6.
@@
$\hskip 56.9055pt+\frac{2\eta}{\sqrt{\lambda}}\frac{\left(\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\right)^{2}+\left(\max_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\right)^{2}}{2}$
@@
where the last inequality holds by Lemma 6.
@@
$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)$ (36)
@@
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}$ $=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}$
$\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (37)
where the last inequality holds by Lemma 6.
@@
$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{16\eta}{\kappa\sqrt{\lambda}}d\log(1+\frac{K}{d\lambda})\right)+\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right)$
$\hskip 28.45274pt+H\sqrt{2KH\log\Big{(}\frac{2}{\delta}\Big{)}}$
$\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1}\big{)}$
@@
Appendix H Proof of Theorem 3

### v3

@@ F.3 Proof of Lemma 3
Lemma 3 follows directly by substituting the confidence set $\widehat{\mathcal{C}}_{k,h}$ defined in Lemma 2, into Lemma 7. This finishes the proof. ∎
F.4 Proof of Theorem 2
@@
=\ = $\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{\mathtt{term~{}(c)}}+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})-p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h}))\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{\mathtt{term~{}(d)}}$
$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{\mathtt{term~{}(e)}}.$
@@
=\ = $\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\Big{\lvert}\Big{(}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime}}-p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\phi_{k,h}^{s^{\prime\prime\prime}}\Big{)}^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\Big{\rvert}\cdot\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
@@
$\hskip 28.45274pt+\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\xi_{k,h})\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\sum_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime\prime}}(\xi_{k,h})\cdot\big{\lvert}(\phi_{k,h}^{s^{\prime\prime\prime}})^{\top}(\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h})\big{\rvert}$
@@
where and the fourth inequality is because by Lemma 15 and the fact ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
$\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}}\leq\big{\lVert}\widetilde{\theta}_{k+1,h}-\widetilde{\theta}_{k,h}\big{\rVert}_{\widetilde{\mathcal{H}}_{k,h}}\leq 2\eta\lVert\nabla\ell_{k,h}(\widetilde{\theta}_{k,h})\rVert_{\widetilde{\mathcal{H}}_{k,h}^{-1}}\leq\frac{2\eta}{\sqrt{\lambda}}\lVert\nabla\ell_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2},$
and since $\nabla\ell_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
$\lVert\nabla\ell_{k,h}(\widetilde{\theta}_{k,h})\rVert_{2}\leq\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}+\Big{\lVert}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\phi_{k,h}^{s^{\prime}}\Big{\rVert}_{2}\leq 2\max_{s^{\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime}}\big{\rVert}_{2}\leq 2.$
@@
$\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\big{\lVert}\bar{\phi}_{k,h}^{s^{\prime}}-\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\frac{8\eta}{\sqrt{\lambda}}\sum_{k=1}^{K}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}$
@@
where the last inequality holds by $\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right)$ in Lemma 6.
@@
$\hskip 56.9055pt+\frac{2\eta}{\sqrt{\lambda}}\left(\Big{(}\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\Big{)}^{2}+\Big{(}\max_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\Big{)}^{2}\right)$
@@
where the last inequality holds by $\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right)$ in Lemma 6.
@@
$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{32\eta}{\kappa\sqrt{\lambda}}d\log\left(1+\frac{K}{d\lambda}\right)\right)$ (26)
@@
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (27)
where the last inequality holds by $\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right)$ in Lemma 6.
@@
$\leq H^{2}\widetilde{\beta}_{K}\left(\sqrt{2dK\log\left(1+\frac{K}{d\lambda}\right)}+\frac{32\eta}{\kappa\sqrt{\lambda}}d\log\left(1+\frac{K}{d\lambda}\right)\right)+\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right)$
$\hskip 28.45274pt+H\sqrt{2KH\log\frac{2}{\delta}}$
$\leq\widetilde{\mathcal{O}}\big{(}dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1}\big{)}.$
@@
Appendix G Omitted Proofs for Section 6
G.1 Proof of Theorem 3